OpenStudy (anonymous):

Find conman solution algebraically and express itrrational roots in simplest radical form. x^2+y^2=24 x+y=8

OpenStudy (shubhamsrg):

where are you stuck ?

OpenStudy (anonymous):

I dont know anything at all to be honest.

OpenStudy (anonymous):

About this topic.

OpenStudy (anonymous):

You could start with thes second equation and find an expression for x= (or y= if you choose) and then substitute that into the first equation.

OpenStudy (anonymous):

I dont get anything =/ second all Just go to shift x to other side right and makes t y=x-8

OpenStudy (anonymous):

y=8-x

OpenStudy (anonymous):

x^2+(8-x)^2=24

OpenStudy (anonymous):

x^2+64-16x+x^2=24

OpenStudy (anonymous):

2x^2-16x+40=0

OpenStudy (anonymous):

x^2-8x+20=0

OpenStudy (anonymous):

so y cnt be x-8 but has to 8-x?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

(x-10)(x+2)=0 x=10 x=-2

OpenStudy (anonymous):

so those will be the solutions right?

OpenStudy (anonymous):

find b's also..

OpenStudy (anonymous):

I meant y

OpenStudy (anonymous):

because these eq. are two variable..

OpenStudy (anonymous):

could you just guide me through like what to do?

OpenStudy (anonymous):

choose one of the eq. it does not matter which one, x+y=8 is easier one than put for x=10 find y for x=-2 find y

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

hold on I made a mistake..

OpenStudy (anonymous):

until this it is true, but I made a mistake when I factored it out.. x^2-8x+20=0

OpenStudy (anonymous):

ax^2+bx+c=0 d=b^2-4ac d=8^2-4*1*20=-16<0 no real solution..

OpenStudy (anonymous):

it means that they are not crossing each other..

OpenStudy (anonymous):

sorry about that..

OpenStudy (anonymous):

root are complex number..

OpenStudy (anonymous):

its okay thanks for your time, really appreciate all this walk through.

OpenStudy (anonymous):

you can see it here

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

np